Freeport mcMoRan projects that in 2010 world grain supply will be 1.8 trillion metric tons and the supply will be only 3/4 of the world demand. What will the world grain demand be in 2010?
Answer by checkley71(8403)

1.8 trillion = 3/4x or x=1.8t(4/3) or x=7.2/3 or x=2.4 trillion